# Traceable Black-Box Watermarks For Federated Learning

### v3

@@ A.11 More Discussions and Future Directions
In our work, we adopt the benchmark FL paradigm, FedAvg, where each local client receives an identical model. However, since $\mathbf{TraMark}$ injects watermarks within the designated watermarking region, malicious clients may collude to identify its location by comparing their identical main task parameters. This limitation can be solved by leveraging personalized FL (Tan et al. 2022; T Dinh, Tran, and Nguyen 2020; Wu, He, and Chen 2020), where the server assigns unique model weights to each client, ensuring distinct watermarked models and enhancing security.
Computational Overhead.
Recall that in our defense model, we assume the server has sufficient computational resources to perform the watermarking process. However, in practical scenarios, computational resources may be limited. Since $\mathbf{TraMark}$ applies watermarking in every training round, this could introduce a non-negligible computational overhead, which increases linearly with the number of participating local clients. Nevertheless, in this work, we focus on cross-silo FL settings, such as collaborations among several hospitals or institutions, where the number of clients is relatively small. In such cases, the server is more likely to have sufficient computational resources, making the additional overhead manageable. Addressing the broader challenge of reducing $\mathbf{TraMark}$’s computational cost in cross-device FL systems remains an open problem for future work.
Theoretical Analysis.
@@
Although $\mathbf{TraMark}$ demonstrates strong empirical performance, there remains a gap in providing a theoretical guarantee for ensuring the traceability of watermarked models. We leave this theoretical analysis as future work.

### v4

@@ A.10 More Discussions and Future Directions
Malicious Client Collusion. In our work, we adopt the benchmark FL paradigm, FedAvg, where each local client receives an identical model. However, since $\mathbf{TraMark}$ injects watermarks within the designated watermarking region, malicious clients may collude to identify its location by comparing their identical main task parameters. This limitation can be solved by leveraging personalized FL (Tan et al., 2022; T Dinh et al., 2020; Wu et al., 2020), where the server assigns unique model weights to each client, ensuring distinct watermarked models and enhancing security.
Computational Overhead with Verification. While we assume the server possesses sufficient computational resources, practical constraints require a careful evaluation of the maximum potential overhead. Having discussed the isolated injection cost in the previous section, we report the cumulative server-side time on CIFAR-10 in Table 9. This metric encompasses model aggregation, watermark injection, and watermark verification. Under this comprehensive metric, $\mathbf{TraMark}$ exhibits a higher total execution time ($74.35$s). However, as emphasized in our defense model, watermark verification is an on-demand audit step performed only when leakage is suspected or at the end of training. Consequently, this value represents a worst-case scenario and should not be interpreted as the standard per-round latency incurred during normal training.
Table 9: Comparison of full server-side computational time (seconds). The reported value includes the time for aggregation, watermark injection, and verification. Note that FedAvg involves only aggregation.
Method FedAvg WAFFLE FedTracker $\mathbf{TraMark}$
Total Time (Agg. + Inject. + Verify.) $1.03$ $12.03$ $13.32$ $74.35$
False Positive Cases. A common challenge faced by black-box watermarking methods is the occurrence of false positives. Specifically, given an unwatermarked model, a black-box watermarking method may still output a prediction, which has a chance of being incorrectly interpreted as watermarked. For our proposed method, $\mathbf{TraMark}$, the expected false positive rate is $1/r$, where $r$ is the number of output labels of the model. Consequently, for large-scale datasets such as Tiny-ImageNet, which contains $200$ classes, the false positive rate is negligible. However, for smaller datasets, the issue becomes more pronounced. Addressing false positives in small-scale datasets is an important direction for future work.
@@
Theoretical Analysis. Although $\mathbf{TraMark}$ demonstrates strong empirical performance, there remains a gap in providing a theoretical guarantee for ensuring the traceability of watermarked models. We leave this theoretical analysis as future work.
A.11 Discussion on Communication Overhead
We analyze the communication overhead of $\mathbf{TraMark}$ in practical network deployments, specifically distinguishing between cross-silo and cross-device settings. In our primary target scenario, cross-silo FL (e.g., collaborations between institutions), systems typically utilize standard TCP/IP protocols where the server communicates with each client via individual point-to-point (unicast) connections. In this context, transmitting $n$ personalized models to $n$ clients consumes the exact same bandwidth as transmitting one identical global model $n$ times, as the total data volume remains unchanged. Since $\mathbf{TraMark}$ modifies only parameter values without altering the model architecture or size, it introduces no additional communication load in cross-silo deployments.
In broadcast-capable environments (e.g., cross-device FL), while distributing distinct models theoretically reduces broadcast efficiency, we argue that this overhead is justifiable for two reasons. First, model differentiation is a fundamental prerequisite for traceability; identifying a specific leaker inherently requires unique model instances, which is a requirement of the traitor tracing problem rather than a limitation specific to $\mathbf{TraMark}$. Second, in settings where multiple devices belong to a single user, per-device watermarking is unnecessary, allowing the same watermarked model to be shared across all devices owned by the same identity. Finally, $\mathbf{TraMark}$ maintains a communication cost profile consistent with existing state-of-the-art traceable methods like FedTracker, which similarly rely on distributing personalized model states.
A.12 Performance Fairness Across Clients
In the main text, we reported the average MA to evaluate the utility of the watermarked models. Here, we provide a detailed breakdown of the per-client accuracy. Specifically, we report the individual testing accuracy of the personalized models for all participating clients (Client ID 099) on both CIFAR-10 and CIFAR-100 datasets. The experiments were conducted with a fixed random seed (0). The results are summarized in Table 10.
Table 10: Per-client MA on CIFAR-10 and CIFAR-100. The results demonstrate high uniformity across all participants.
Dataset ID: 0 ID: 1 ID: 2 ID: 3 ID: 4 ID: 5 ID: 6 ID: 7 ID: 8 ID: 9
CIFAR-10 88.48 88.69 88.33 88.36 88.08 88.40 88.13 88.30 88.37 88.14
CIFAR-100 61.24 61.40 60.79 60.85 61.22 61.06 60.89 61.35 60.99 61.12
As shown in the table, the performance disparity among clients is negligible. On the CIFAR-10 dataset, the standard deviation of accuracy across clients is approximately $0.17\%$, while on the more complex CIFAR-100 dataset, it remains low at approximately $0.19\%$. We observed no outlier cases where a specific client received a poorly performing model. These results confirm that $\mathbf{TraMark}$ preserves equitable model utility for all participants in the federated system.
A.13 Discussion on the Trade-off between Privacy and Traceability
A key motivation for using FL is preserving data privacy. Adding traceability raises questions about whether these guarantees still hold. Here, we discuss this trade-off in $\mathbf{TraMark}$ and show that although model-level anonymity is reduced, data privacy remains fully preserved.
In standard FL systems such as FedAvg, privacy comes from keeping raw data local while sharing only model updates. $\mathbf{TraMark}$ follows the same process. The server already receives separate updates from each client, and our method does not change this communication protocol or access any local data. The traceability mechanism is based only on OOD triggers generated by the server, which do not reveal information about private training data.
The trade-off introduced by $\mathbf{TraMark}$ concerns model anonymity. In settings that use secure aggregation, the server observes only the aggregated update and cannot identify individual contributions. In contrast, $\mathbf{TraMark}$ requires distributing and verifying personalized models, which means the server knows which model belongs to which client. This reduction in anonymity is a functional requirement for accountability. In many cross-silo FL applications, the server is the model owner and must be able to identify model leakage to protect intellectual property. In such cases, full anonymity would prevent accountability and allow model leakers to go untraceable.
A.14 Use of LLM Statement
We used LLM solely for grammar checking and polishing the writing of this manuscript.